According to a research institution, men spent an average of $136.89 on Valentines Day gifts in 2009. Assume the standard deviation for this population is $40 and that it is normally distributed. A random sample of 10 men who celebrate Valentines Day was selected. Complete parts a through e. a. Calculate the standard error of the mean. Round to two decimal places as needed.) b. What is the probability that the sample mean will be less than $130? P (x<$130) Round to four decimal places as needed.) c. What is the probability that the sample mean will be more than $1451? P (->$145)- ] Round to four decimal places as needed.) d. What is the probability that the sample mean will be between $115 and $155? P (S115 sxsS155) Round to four decimal places as needed) e. Identify the symmetrical interval that includes 95% of the sample means if the true population mean is S 136.89 Round to the nearest dollar as needed.)
Answer #1

Solution:-

a) Standard error of the mean is 12.649.

b) The probability that the sample mean will be less than 130 is 0.293.

x = 130

By applying normal distruibution:-

z = \frac{x-\mu }{\frac{\sigma }{\sqrt{n}}}

z = - 0.545

P(z < - 0.545) = 0.293.

c) The probability that the sample mean will be more than 145 is 0.005.

x = 145

By applying normal distruibution:-

z = \frac{x-\mu }{\frac{\sigma }{\sqrt{n}}}

z = 2.565

P(z > 2.565) = 0.005.

d) The probability that the sample mean will be between $115 and $155 is 0.882.

x1 = 115

x2 = 155

By applying normal distruibution:-

z = \frac{x-\mu }{\frac{\sigma }{\sqrt{n}}}

z1 = - 1.73

z2 = 1.43

P(- 1.73 < z < 1.43) = P(z > - 1.73) - P(z > 1.43)

P(- 1.73 < z < 1.43) = 0.958 - 0.076

P(- 1.73 < z < 1.43) = 0.882